# Creating Multi-Level Skill Hierarchies in Reinforcement Learning

### v1

@@ Abstract
What is a useful skill hierarchy for an autonomous agent? We propose an answer based on the graphical structure of an agent’s interaction with its environment. Our approach uses hierarchical graph partitioning to expose the structure of the graph at varying timescales, producing a skill hierarchy with multiple levels of abstraction. At each level of the hierarchy, skills move the agent between regions of the state space that are well connected within themselves but weakly connected to each other. We illustrate the utility of the proposed skill hierarchy in a wide variety of domains in the context of reinforcement learning.
@@ 1 Introduction
How can an agent autonomously develop an action hierarchy as it interacts with its environment? This is a fundamental open question in artificial intelligence. Before answering this algorithmic question, it is useful to first ask a conceptual question: What is a useful action hierarchy? Here we focus on this conceptual question to provide a useful foundation for future algorithmic development.
We propose a characterisation of a useful action hierarchy based on a graphical representation of an agent’s interaction with its environment. We first partition this graph at multiple scales, exposing the structure of the environment at various levels of granularity, then define actions that efficiently move an agent between neighbouring clusters in the partitions. The outcome is an action hierarchy that enables the agent to efficiently interact with its environment at multiple time scales. In a diverse set of environments, the proposed characterisation translates into action hierarchies that are intuitively appealing, improve learning performance, and has desirable scaling properties.
Our approach has two key characteristics. First, we partition the interaction graph by maximising modularity, a measure of partition quality that generates clusters that are strongly connected within themselves but weakly connected to each other. Secondly, our approach produces a multi-level hierarchy, where lower-level actions are naturally composed into higher-level actions. Multi-level action hierarchies naturally support the ability to act, plan, explore, and learn over varying timescales, offering many concrete benefits over unstructured collections of actions that are not organised into a hierarchy.
@@ 2 Background
We use the reinforcement learning framework, modelling an agent’s interaction with its environment as a finite Markov Decision Process (MDP). An MDP is a six-tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{D},\gamma)$, where $\mathcal{S}$ is a finite set of states, $\mathcal{A}$ is a finite set of actions, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is a transition function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is a reward function, $\mathcal{D}:\mathcal{S}\rightarrow[0,1]$ is an initial state distribution, and $\gamma\in[0,1]$ is a discount factor. Let $\mathcal{A}(s)$ denote the set of actions available in state $s\in\mathcal{S}$. At decision stage $t$, $t\geq 0$, the agent observes state $s_{t}\in\mathcal{S}$ and executes action $a_{t}\in\mathcal{A}(s_{t})$. Consequently, at decision stage $t+1$, the agent receives a numerical reward, $r_{t+1}\in\mathcal{R}$, and observes the next state, $s_{t+1}\in\mathcal{S}$. The return at decision stage $t$ is the discounted sum of future rewards, $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}$. A policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is a mapping from state-action pairs to probabilities. The agent’s objective is to learn a policy that maximises the expected return. The state-transition graph of an MDP is a weighted, directed graph whose nodes represent the states of the MDP and whose edges represent possible transitions between these states via primitive actions. An edge $(u,v)$ exists on the graph if it is possible to transition from state $u\in\mathcal{S}$ to state $v\in\mathcal{S}$ by taking some action $a\in\mathcal{A}(u)$. Unless stated otherwise, we use uniform edge weights of $1$.
@@
The actions of an MDP take exactly one decision stage to execute; we refer to them as primitive actions. Using primitive actions, it is possible to define abstract actions, or skills, whose execution can take a variable number of decision stages. Primitive and abstract actions can be combined to form complex action hierarchies. We represent skills using the options framework Sutton et al. (1999a). An option $o$ is a three-tuple $(\mathcal{I}_{o},\pi_{o},\beta_{o})$, where $\mathcal{I}_{o}\subset\mathcal{S}$ is the initiation set, specifying the set of states in which the option can start execution, $\pi_{o}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the option policy, and $\beta_{o}:\mathcal{S}\rightarrow[0,1]$ is the termination condition, specifying the probability of option termination in a given state. An option policy is ultimately defined over primitive actions—because these are the fundamental units of interaction between the agent and its environment—but this can be done indirectly by allowing options to call other options, making it possible for agents to operate with hierarchies of primitive and abstract actions.
@@ 3 Related Work
Although many existing approaches to skill discovery use the state-transition graph, all such approaches produce skill hierarchies with only a single level of skills above primitive actions. Multi-level skill hierarchies are essential when solving complex tasks, which often require agents to act, plan, explore, and learn over varying timescales. They offer many concrete benefits to reinforcement learning agents. First, when using a multi-level hierarchy, an agent can learn about not only the skill it is currently executing but also about any lower-level skills it calls upon. Secondly, arranging skills into multi-level hierarchies allows them to be updated in a modular fashion. For instance, any improvements to the internal policy of a lower-level skill would be immediately reflected in all higher-level skills that call it. Thirdly, the ability to form multi-level skill hierarchies supports a continual learning process, where an agent combines existing skills to produce new skills in an open-ended manner.
While existing graph-based methods do not learn multi-level hierarchies, policy-gradient methods have made some progress towards this goal. Bacon et al. (2017) extended policy-gradient theorems Sutton et al. (1999b) to allow the learning of option policies and termination conditions in a two-level hierarchy. Riemer et al. Riemer et al. (2018) further generalised these theorems to support multi-level hierarchies. Fox et al. (2017) propose an imitation learning method that finds the multi-level skill hierarchy most likely to generate a given set of example trajectories. Levy et al. (2019) propose a method for learning multi-level hierarchies of goal-directed policies, with each level of the hierarchy producing a subgoal for the lower-levels to navigate towards. However, these methods are not without their limitations. Unlike the approach proposed here, they all require the number of hierarchy levels to be pre-defined instead of finding a suitable number automatically. They also make simplifying assumptions, such as that all skills are available in all states, and target different types of problems than we do, such as imitation-learning or goal-directed problems with high-dimensional or continuous state-spaces.
Our approach is most directly related to skill discovery methods that use graph partitioning  (Mannor et al., 2004; Davoodabadi and Beigy, 2011; Metzen, 2013; Shoeleh and Asadpour, 2017; Xu et al., 2018; Farahani and Mozayani, 2019). Three of these methods use the concept of modularity, which is also central to our approach. One such approach is to generate a series of possible partitions by successively removing the edge with the highest edge betweenness from a graph, then selecting the partition with the highest modularity (Davoodabadi and Beigy, 2011). A second approach is to generate a partition using the label propagation algorithm and then to merge neighbouring clusters until no gain in modularity is possible (Farahani and Mozayani, 2019). In these two approaches, the final partition will maximise modularity with respect to only the initial clusters identified by other methods; in other words, the final partition will generally not be the one that maximises modularity overall. The label propagation method runs in near-linear time, whereas the edge betweenness method has a time complexity of $O(m^{2}n)$ on a graph with $m$ edges and $n$ nodes. A third approach is by Xu et al. (2018), who use the Louvain algorithm to find a partition that maximises modularity but, unlike our approach, define skills only for moving between clusters in the highest-level partition, discarding all lower-level partitions. Using only the highest-level partition produces skills for navigating the state-space at only a high level, whereas our approach produces skills for navigating the state-space at varying timescales by using the full cluster hierarchy.
@@
Another approach to skill discovery is to identify useful subgoals and define skills for navigating to them. Suggestions have often been inspired by the concept of “bottleneck” states. They include states that are on the border of strongly-connected regions of the state-space (Menache et al., 2002), states that allow transitions between different regions of the state-space (Şimşek and Barto, 2004), and states that lie on the shortest path between many other pairs of states Şimşek and Barto (2009). To identify such states, several approaches use graph centrality measures (Şimşek and Barto, 2009; Moradi et al., 2010; Rad et al., 2010; Imanian and Moradi, 2011; Moradi et al., 2012). Others use graph partitioning algorithms to identify meaningful regions of the state-space, then identify subgoals as the nodes bordering them (Metzen, 2013; Menache et al., 2002; Şimşek et al., 2005; Kazemitabar and Beigy, 2009; Entezari et al., 2010; Bacon and Precup, 2013; Taghizadeh and Beigy, 2013; Kazemitabar et al., 2018). The bottleneck concept has also inspired non-graphical approaches to skill discovery, such as seeking to identify states that are visited frequently on successful trajectories but infrequently on unsuccessful ones (McGovern and Barto, 2001). Alternatively, it has been proposed that “landmark” states found at the centre of strongly-connected regions of the state-space can be used as subgoals Ramesh et al. (2019).
@@
Several approaches have used the graph Laplacian (Machado et al., 2017; Jinnai et al., 2019) to identify skills that are specifically useful for efficiently exploring the state space. These methods produce skills that aim to minimise the expected number of actions required to navigate between any two states, allowing efficient navigation between distant areas of the state space. It is unclear how to arrange such skills to form multi-level skill hierarchies. In contrast, we propose a principled approach to characterising skills that are both useful and naturally form a multi-level hierarchy.
4 Proposed Approach
We identify partitions of the state-transition graph that maximise modularity (Newman and Girvan, 2004; Leicht and Newman, 2008). A partition of a graph is a division of its nodes into mutually exclusive groups, called clusters. Given a set of clusters $C=\{c_{1},c_{2},\ldots,c_{k}\}$ forming a partition of a graph, the modularity of the partition is
$\sum_{i=1}^{k}e_{ii}-\rho a^{2}_{i}\text{ ,}$
where $e_{ii}$ denotes the proportion of total edge weight in the graph that connects two nodes in cluster $c_{i}$, and $a_{i}$ denotes the proportion of total edge weight in the graph with at least one end connected to a node in cluster $c_{i}$. A resolution parameter $\rho$ controls the relative importance of $e_{ii}$ and $a_{i}$. Intra-cluster edges contribute to both $e_{ii}$ and $a_{i}$ while inter-cluster edges contribute only to $a_{i}$. So, as $\rho$ increases, large clusters with many inter-cluster edges are penalised to a higher degree, leading to partitions with smaller clusters.
A partition that maximises modularity will have dense connections within its clusters but sparse connections between them. In the context of reinforcement learning, these clusters correspond to regions of the state space that are easy to navigate within but difficult to navigate between. The proposed skill hierarchy gives an agent the ability to efficiently move between such regions.
Finding a partition that maximises modularity for a given graph is NP-complete (Brandes et al., 2006). Therefore, when working with large graphs, approximation algorithms are needed. The most widely used approximation algorithm is the Louvain algorithm (Blondel et al., 2008), an agglomerative hierarchical graph clustering algorithm. While no formal analysis exists, the runtime of the Louvain algorithm has been observed empirically to be linear in the number of graph edges (Lancichinetti and Fortunato, 2009).
The Louvain algorithm starts by placing each node of the graph in its own cluster. Nodes are then iteratively moved locally, from their current cluster to a neighbouring cluster, until no gain in modularity is possible. This results in a revised partition corresponding to a local maximum of modularity with respect to local node movement. This revised partition is used to define an aggregate graph as follows: each cluster in the partition is represented as a single node in the aggregate graph, and a directed edge is added to the aggregate graph if there is at least one edge that connects neighbouring clusters in that direction. This process is then repeated on the aggregate graph, and then on the next aggregate graph, and so on, until an iteration is reached with no modularity gain. For more details please refer to Blondel et al. (2008)Arenas et al. (2007), and the pseudocode in Appendix B.
The algorithm’s output is a series of partitions of the input graph. This series of partitions has a useful structure: multiple clusters found in one partition are merged into a single cluster in the next partition. In other words, the output is a hierarchy of clusters, with earlier partitions containing many smaller clusters which are merged into fewer larger clusters in later partitions. This hierarchical structure forms the basis of the multi-level skill hierarchy we propose for reinforcement learning.
Let $h$ denote the number of partitions returned by the algorithm. We use each of the $h$ partitions to define a single layer of skills, resulting in a hierarchy with $h$ levels above primitive actions. Each level of the hierarchy contains one or more skills for efficiently navigating between neighbouring clusters. Specifically, we define an option for navigating from a cluster $c_{i}$ to a neighbouring cluster $c_{j}$ as follows: the initiation set consists of all states in $c_{i}$; the option policy efficiently takes the agent from a given state in $c_{i}$ to a state in $c_{j}$; the option terminates with probability $1$ in states in $c_{j}$, with probability 00 otherwise.
Taking advantage of the natural hierarchical structure of the partitions produced by the Louvain algorithm, we compose the skills at one level of the hierarchy to define the skills at the next level. That is, at each level of the hierarchy, option policies will be defined over only the actions (options or primitive actions) in the level below, with options at only the first level of the hierarchy using policies defined over primitive actions. We call the resulting set of skills the Louvain skill hierarchy.

### v2

@@ Abstract
What is a useful skill hierarchy for an autonomous agent? We propose an answer based on a graphical representation of how the interaction between an agent and its environment may unfold. Our approach uses modularity maximisation as a central organising principle to expose the structure of the interaction graph at multiple levels of abstraction. The result is a collection of skills that operate at varying time scales, organised into a hierarchy, where skills that operate over longer time scales are composed of skills that operate over shorter time scales. The entire skill hierarchy is generated automatically, with no human intervention, including the skills themselves (their behaviour, when they can be called, and when they terminate) as well as the hierarchical dependency structure between them. In a wide range of environments, this approach generates skill hierarchies that are intuitively appealing and that considerably improve the learning performance of the agent.
@@ 1 Introduction
One of the most important open problems in artificial intelligence is how to make it possible for autonomous agents to develop useful action hierarchies on their own, without any input from humans, including system designers and domain specialists. Before addressing this algorithmic question, it is useful to first consider a conceptual one: What constitutes a useful action hierarchy? Here we focus on this conceptual question with the aim of providing a useful foundation for algorithmic development.
Our primary contribution is a characterisation of a useful action hierarchy. In defining this action hierarchy, we use no information other than a graphical representation of how the interaction between the agent and its environment may unfold. When this information is not known a priori, it would be discovered naturally by the agent as it operates in its environment. Beyond this interaction graph, no particular domain knowledge is needed. Hence, our approach is applicable broadly.
Multiple strands of earlier work have used the interaction graph as a basis for defining collections of actions. The main novelty in our approach is our use of modularity maximisation as a central organising principle to expose the structure of the interaction graph at multiple levels of abstraction. The outcome is an action hierarchy that enables the agent to explore its environment efficiently at multiple time scales.
Our approach yields an action hierarchy with four desirable properties. First, it contains actions that operate at a wide range of time scales. This is necessary to solve complex problems, which require agents to be able to act, learn, and plan at varying time scales. Secondly, the actions are naturally organised into a hierarchy, with actions that operate over longer time scales being composed of actions that operate over shorter time scales. This hierarchical structure offers substantial benefits over unstructured collections of actions. For example, it allows an agent to learn about not only the action it is currently executing but also all lower-level actions that are called in the process. In addition, a hierarchical structure allows actions to be updated in a modular fashion. For instance, any improvements to an action would be immediately reflected in all higher-level actions that call it. Thirdly, the action hierarchy is fully specified. This includes when each action can be selected for execution, how exactly it behaves, and when it terminates. It also includes the number of levels in the hierarchy and the exact dependency structure between the actions. Fourthly, the action hierarchy is generated automatically, with no human intervention.
In a diverse set of environments, the proposed approach translates into action hierarchies that are intuitively appealing. When evaluated within the context of reinforcement learning, they substantially improve learning performance compared to alternative approaches, with the largest performance improvement observed in the largest environment tested.
An important question for future research is how such an action hierarchy may be learned when the agent has no prior knowledge of the dynamics of the environment. We present an initial exploration of how this may be achieved, with positive results.
@@ 2 Background
We use the reinforcement learning framework, modelling an agent’s interaction with its environment as a finite Markov Decision Process (MDP). An MDP is a six-tuple $(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{D},\gamma)$, where $\mathcal{S}$ is a finite set of states, $\mathcal{A}$ is a finite set of actions, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ is a transition function, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is a reward function, $\mathcal{D}:\mathcal{S}\rightarrow[0,1]$ is an initial state distribution, and $\gamma\in[0,1]$ is a discount factor. Let $\mathcal{A}(s)$ denote the set of actions available in state $s\in\mathcal{S}$. At decision stage $t$, $t\geq 0$, the agent observes state $s_{t}\in\mathcal{S}$ and executes action $a_{t}\in\mathcal{A}(s_{t})$. Consequently, at decision stage $t+1$, the agent receives reward $r_{t+1}\in\mathbb{R}$ and observes new state $s_{t+1}\in\mathcal{S}$. The return at decision stage $t$, denoted by $G_{t}$, is the discounted sum of future rewards, $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}$. A policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ specifies the probability of selecting action $a\in\mathcal{A}$ in state $s\in\mathcal{S}$. The objective is to learn a policy that maximises the expected return.
The state transition graph of an MDP is a weighted, directed graph whose nodes correspond to the states of the MDP and whose edges correspond to possible transitions between states. Specifically, an edge $(u,v)$ exists on the graph if it is possible to transition from state $u\in\mathcal{S}$ to state $v\in\mathcal{S}$ by taking some action $a\in\mathcal{A}(u)$. In this paper, we use uniform edge weights of $1$.
@@
The actions of an MDP take exactly one decision stage to execute. We refer to them as primitive actions. Using primitive actions, it is possible to define abstract actions, also known as skills, whose execution can take a variable number of decision stages. Furthermore, primitive and abstract actions can be combined to form complex action hierarchies. In this work, we represent abstract actions using the options framework Sutton et al. (1999a); Precup (2000). An option $o$ is a three-tuple $(\mathcal{I}_{o},\pi_{o},\beta_{o})$, where $\mathcal{I}_{o}\subset\mathcal{S}$ is the initiation set, specifying the set of states in which the option can start execution, $\pi_{o}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ is the option policy, and $\beta_{o}:\mathcal{S}\rightarrow[0,1]$ is the termination condition, specifying the probability of option termination in a given state. An option policy is ultimately defined in terms of primitive actions—because primitive actions are the fundamental units of interaction between the agent and its environment—but this can be done indirectly by allowing options to call other options, making it possible for agents to act, learn, and plan with hierarchies of primitive and abstract actions.
3 Proposed Approach
To define a skill hierarchy, we use modularity maximisation as a central organising principle, applied at multiple time scales. Specifically, we represent the possibilities of interaction between the agent and its environment as a graph and identify partitions of this graph that maximise modularity (Newman and Girvan, 2004; Leicht and Newman, 2008; Arenas et al., 2007).
A partition of a graph is a division of its nodes into mutually exclusive groups, called clusters. The modularity of a partition composed of a set of clusters $C=\{c_{1},c_{2},\ldots,c_{k}\}$ is
$\sum_{i=1}^{k}e_{ii}-\rho a^{2}_{i}\text{ ,}$
where $e_{ii}$ denotes the proportion of total edge weight in the graph that connects two nodes in cluster $c_{i}$, and $a_{i}$ denotes the proportion of total edge weight in the graph with at least one end connected to a node in cluster $c_{i}$. A resolution parameter $\rho>0$ controls the relative importance of $e_{ii}$ and $a_{i}$. Intra-cluster edges contribute to both $e_{ii}$ and $a_{i}$ while inter-cluster edges contribute only to $a_{i}$. A partition that maximises modularity will have relatively dense connections within its clusters and relatively sparse connections between its clusters.
Finding a partition that maximises modularity for a given graph is NP-complete (Brandes et al., 2006). Therefore, when working with large graphs, approximation algorithms are needed. The most widely used approximation algorithm is the Louvain algorithm (Blondel et al., 2008), which is an agglomerative hierarchical graph clustering approach. While no formal analysis exists, the runtime of the Louvain algorithm has been observed empirically to be linear in the number of graph edges (Lancichinetti and Fortunato, 2009). It has been successfully applied to graphs with millions of nodes and billions of edges Blondel et al. (2008).
An important feature of the Louvain algorithm is that, as a hierarchical graph clustering method, it exposes the structure of a graph at multiple levels of granularity. Specifically, the output of the Louvain algorithm is a sequence of partitions of the input graph. This sequence has a useful structure: multiple clusters found in one partition in the sequence are merged into a single cluster in the next partition in the sequence. In other words, the output is a hierarchy of clusters, with earlier partitions containing many smaller clusters that are merged into fewer larger clusters in later partitions. This hierarchical structure forms the basis of our characterisation of a useful multi-level skill hierarchy.
The Louvain algorithm starts by placing each node of the graph in its own cluster. Nodes are then iteratively moved locally, from their current cluster to a neighbouring cluster, until no gain in modularity is possible. This results in a revised partition corresponding to a local maximum of modularity with respect to local node movement. The revised partition is used to define an aggregate graph as follows: each cluster in the partition is represented as a single node in the aggregate graph, and a directed edge is added to the aggregate graph if there is at least one edge that connects neighbouring clusters in the corresponding direction. This process is then repeated on the aggregate graph, and then on the next aggregate graph, and so on, until an iteration is reached with no modularity gain. Pseudocode for the Louvain algorithm is presented in Section H of the supplementary material.
Let $h$ denote the number of partitions returned by the Louvain algorithm when applied to the state transition graph. We use each of the $h$ partitions to define a single layer of skills, resulting in an action hierarchy with $h$ levels of abstract actions (skills) above primitive actions. Each level of the hierarchy contains one or more skills for efficiently navigating between neighbouring clusters of the state transition graph. Specifically, we define an option for navigating from a cluster $c_{i}$ to a neighbouring cluster $c_{j}$ as follows: the initiation set consists of all states in $c_{i}$; the option policy efficiently takes the agent from a given state in $c_{i}$ to a state in $c_{j}$; the option terminates with probability $1$ in states in $c_{j}$, with probability 00 otherwise.
Taking advantage of the natural hierarchical structure of the partitions produced by the Louvain algorithm, we compose the skills at one level of the hierarchy to define the skills at the next level up. That is, at each level of the hierarchy, option policies call actions (options or primitive actions) from the level below, with primitive actions being called directly by option policies from only the first level of the hierarchy. We call the resulting set of skills the Louvain skill hierarchy.
@@ 4 Related Work
There have been earlier approaches to skill discovery using the state transition graph. The approach we propose here differs from them in two fundamental ways. First, it is novel in its use of modularity maximisation as a central organising principle for skill discovery. Secondly, it produces a multi-level hierarchy, whereas existing graph-based approaches produce hierarchies with only a single level of skills above primitive actions.
@@
Many existing approaches to skill discovery use the state transition graph to identify useful subgoal states and define skills that efficiently take the agent to these subgoals. Suggestions for useful subgoals have often been inspired by the concept of a “bottleneck”. They include states that are on the border of strongly-connected regions of the state space (Menache et al., 2002), states that allow transitions between different regions of the state space (Şimşek and Barto, 2004), and states that lie on the shortest path between many pairs of states Şimşek and Barto (2009). To identify such states, several approaches use graph centrality measures (Şimşek and Barto, 2009; Moradi et al., 2010; Rad et al., 2010; Imanian and Moradi, 2011; Moradi et al., 2012). Others use graph partitioning algorithms (Menache et al., 2002; Metzen, 2013; Şimşek et al., 2005; Kazemitabar and Beigy, 2009; Entezari et al., 2010; Bacon and Precup, 2013; Taghizadeh and Beigy, 2013; Kazemitabar et al., 2018). Alternatively, it has been proposed that “landmark” states found at the centre of strongly-connected regions of the state-space can be used as subgoals Ramesh et al. (2019).
The proposed approach is most directly related to skill discovery methods that make use of graph partitioning to identify meaningful regions of the state transition graph and define skills for navigating between them (Metzen, 2013; Mannor et al., 2004; Davoodabadi and Beigy, 2011; Shoeleh and Asadpour, 2017; Xu et al., 2018; Farahani and Mozayani, 2019). Three of these methods use the concept of modularity. One such approach is to generate a series of possible partitions by successively removing the edge with the highest edge betweenness from a graph, then selecting the partition with the highest modularity (Davoodabadi and Beigy, 2011). A second approach is to generate a partition using the label propagation algorithm and then to merge neighbouring clusters until no gain in modularity is possible (Farahani and Mozayani, 2019). In these two approaches, modularity maximisation is applied after first producing candidate partitions using different criteria. Consequently, the final partition does not maximise modularity overall. The time complexity of the label propagation method is near-linear in the number of graph edges, whereas the edge betweenness method has a time complexity of $O(m^{2}n)$ on a graph with $m$ edges and $n$ nodes. A third approach is by Xu et al. (2018), who use the Louvain algorithm to find a partition that maximises modularity but, unlike our approach, define skills only for moving between clusters in the highest-level partition, discarding all lower-level partitions. All three methods produce a single level of skills above primitive actions.
@@
Several approaches have used the graph Laplacian (Machado et al., 2017; Jinnai et al., 2019) to identify skills that are specifically useful for efficiently exploring the state space. It is unclear how to arrange such skills to form multi-level skill hierarchies. In contrast, the proposed approach produces a set of skills that are naturally arranged into a multi-level hierarchy.
While existing graph-based methods do not learn multi-level hierarchies, policy-gradient methods have made some progress towards this goal. Bacon et al. (2017) extended policy-gradient theorems Sutton et al. (1999b) to allow the learning of option policies and termination conditions in a hierarchy with a single level of skills above primitive actions. Riemer et al. Riemer et al. (2018) further generalised these theorems to support multi-level hierarchies. Fox et al. (2017) propose an imitation learning method that finds the multi-level skill hierarchy most likely to generate a given set of example trajectories. Levy et al. (2019) propose a method for learning multi-level hierarchies of goal-directed policies, with each level of the hierarchy producing a subgoal for the lower levels to navigate towards. However, these methods are not without their limitations. Unlike the approach proposed here, they all require the number of hierarchy levels to be pre-defined instead of finding a suitable number automatically. They do not judiciously define initiation sets, instead making all skills available in all states. They also target different types of problems than we do, such as imitation-learning or goal-directed problems.